OpenStudy (anonymous):

will medal A. 2 B. 10 C. 9 D. 5

OpenStudy (kaylardz0405):

the question?

Nnesha (nnesha):

The answer is E.

Nnesha (nnesha):

any equation or expression ? what about those options ?

OpenStudy (anonymous):

what is the value of the expression when a=5,b=4,and c=6

OpenStudy (anonymous):



OpenStudy (kendricklamar2014):

Substitute!

OpenStudy (kaylardz0405):

so first multiply 4*5 because a = 5

OpenStudy (anonymous):

20

OpenStudy (kendricklamar2014):

\[\left(\begin{matrix}4 \times 5 \\ 2 \times 4-6\end{matrix}\right)\]

OpenStudy (kaylardz0405):

yes so now multiply 2*4

OpenStudy (anonymous):

8